# Detection and Recovery of Hidden Submatrices

### v1

@@ 1 Introduction
We use standard asymptotic notation. For two positive sequences $\{a_{n}\}$ and $\{b_{n}\}$, we write $a_{n}=O(b_{n})$ if $a_{n}\leq Cb_{n}$, for some absolute constant $C$ and for all $n$; $a_{n}=\Omega(b_{n})$, if $b_{n}=O(a_{n})$; $a_{n}=\Theta(b_{n})$, if $a_{n}=O(b_{n})$ and $a_{n}=\Omega(b_{n})$, $a_{n}=o(b_{n})$ or $b_{n}=\omega(a_{n})$, if $a_{n}/b_{n}\to 0$, as $n\to\infty$. Finally, for $a,b\in\mathbb{R}$, we let $a\vee b\triangleq\max\{a,b\}$ and $a\wedge b\triangleq\min\{a,b\}$. Throughout the paper, $C$ refers to any constant independent of the parameters of the problem at hand and will be reused for different constants. The notation $\ll$ refers to polynomially less than in $n$, namely, $a_{n}\ll b_{n}$ if $\liminf_{n\to\infty}\log_{n}a_{n}<\liminf_{n\to\infty}\log_{n}b_{n}$, e.g., $n\ll n^{2}$, but $n\not\ll n\log_{2}n$. For $n\in\mathbb{N}$, we let $[n]=\{1,2,\dots,n\}$. For a subset $S\subseteq\mathbb{R}$, we let $\mathbbm{1}\{S\}$ denote the indicator function of the set $S$.
@@ 3.2 The recovery problem
Before we present this algorithm, we make a simplifying assumption on the possible set of planted submatrices and then explain how this assumption can be removed. We assume that there is no pair of adjacent submatrices in the underlying planted submatrices $\mathsf{K}^{\star}$. In other words, there is at least one column or row separating any pair of submatrices. Note that in a random ensemble, where the planted submatrices are chosen uniformly at random from ${\cal K}_{k,m,n}^{\mathsf{con}}$, as we show in Subsection 4.5.1, the event of a pair of adjacent submatrices is of small probability, as long as $mk=o(n)$. Nonetheless, below we discuss how this technical assumption can be dropped easily. Similar assumptions are frequently taken when analyzing statistical models inspired by cryo-EM, see, for example [6].
@@
Consider the recovery problem in Definition 3. Then, we have the following set of bounds:
@@
    (Exact Recovery) Consider the peeling estimator in (20). If
    $\liminf_{n\to\infty}\frac{\lambda}{\sqrt{6k^{-1}\log n}}>1,$ (22)
@@
    (Correlated Recovery) Consider the peeling estimator in (20). If
    $\liminf_{n\to\infty}\frac{\lambda}{2\sqrt{2k^{-2}\log n}}>1,$ (23)
@@ 4.5 Proof of Theorem 5
Before delving into the proof, we prove that over the random ensemble, where $\mathsf{K}^{\star}$ is drawn uniformly at random from ${\cal K}_{k,m,n}^{\mathsf{con}}$, the assumption that any two submatrices in $\mathsf{K}^{\star}$ are not adjacent holds with high probability. Thus, if this random ensemble is considered we can safely condition on this high-probability event, and assume that the assumption holds.
4.5.1 Random ensemble
Let ${\cal A},{\cal B}\subset[n]$ be two disjoint sets of $k$ consecutive elements. Let $a_{\max}$ ($b_{\max}$) and $a_{\min}$ ($b_{\min}$) be the largest and smallest elements of ${\cal A}$ (${\cal B}$), respectively. Assume that ${\cal A}\preceq{\cal B}$, in the sense that $a_{\max}\leq b_{\min}$. Then, the ${\cal A}$ and ${\cal B}$ are adjacent if $b_{\min}-a_{\max}=1$. For example, for $k=4$, ${\cal A}=(1,2,3,4)$ and ${\cal B}=(5,6,7,8)$ are adjacent, while ${\cal A}=(1,2,3,4)$ and ${\cal B}=(9,10,11,12)$ are not. Now, let ${\cal E}$ denote the event that for $\mathsf{K}\sim\mathsf{Unif}({\cal K}_{k,m,n}^{\mathsf{con}})$, there exist at least two submatrices in $\mathsf{K}$, say, $\mathsf{S}\times\mathsf{T}$ and $\mathsf{S}^{\prime}\times\mathsf{T}^{\prime}$, such that:
  1. 1.
    Either or both of $\mathsf{S}\preceq\mathsf{S^{\prime}}$ and $\mathsf{T}\preceq\mathsf{T^{\prime}}$ hold.
  2. 2.
    Either or both $(\mathsf{S},\mathsf{S}^{\prime})$ and $(\mathsf{T},\mathsf{T}^{\prime})$ are adjacent.
It is clear that the number of pairs $\mathsf{S}\times\mathsf{T}$ and $\mathsf{S}^{\prime}\times\mathsf{T}^{\prime}$ such that items (1) and (2) above hold is at most $\Theta(n^{2}k)$; the factor of $n^{2}$ reflects the maximum number of choices for $\mathsf{S}\times\mathsf{T}$ and the factor of $k$ reflects the number of ways to have $\mathsf{S}^{\prime}\times\mathsf{T}^{\prime}$ adjacent to $\mathsf{S}\times\mathsf{T}$. Since other submatrices in $\mathsf{K}$ are not necessarily constrained, the total number of matrices $\mathsf{K}$ such that there exist at least two submatrices in $\mathsf{K}$ is at most $\Theta(mkn^{2})\cdot(n-k)\prod_{\ell=2}^{m-1}(n-\ell k)^{2}$, where the factor of $m$ reflects the number of possible pairs of (consecutive) matrices, and $(n-k)\prod_{\ell=2}^{m-1}(n-\ell k)^{2}=\prod_{\ell=1}^{m-1}(n-\ell k)\prod_{\ell=2}^{m-1}(n-\ell k)$ is the number of ways to pick the other submatrices in $\mathsf{K}$, given the first fixed adjacent matrices. Thus,
$\mathbb{P}({\cal E})$ $\leq\frac{\Theta(mkn^{2})\cdot(n-k)\prod_{\ell=2}^{m-1}(n-\ell k)^{2}}{|{\cal K}_{k,m,n}^{\mathsf{con}}|}$ (101)
$=\frac{\Theta(mkn^{2})\cdot(n-k)\prod_{\ell=2}^{m-1}(n-\ell k)^{2}}{\prod_{\ell=0}^{m-1}(n-\ell k)^{2}}$ (102)
$=\frac{\Theta(mk)}{(n-k)}=o(1),$ (103)
assuming that $mk=o(n)$.
@@ 4.5.2 Exact recovery using the peeling estimator
where ${\cal S}(\mathsf{K})\triangleq\sum_{(i,j)\in\mathsf{K}}\mathsf{X}_{ij}$. We next prove the conditions for which $\hat{\mathsf{K}}_{1}(\mathsf{X})=\mathsf{K}^{\star}_{1}$, with high probability, where $\mathsf{K}^{\star}_{1}$ is one out of the $m$ planted submatrices in $\mathsf{K}^{\star}$. To prove the theorem, it suffices to show that ${\cal S}(\mathsf{K}^{\star}_{1})>{\cal S}(\mathsf{K})$for all feasible $\mathsf{K}$ with $\mathsf{K}\neq\mathsf{K}_{1}^{\star}$. Let $\mathsf{D}(\mathsf{K})\triangleq{\cal S}(\mathsf{K}^{\star}_{1})-{\cal S}(\mathsf{K})$. Note that
$\mathsf{D}(\mathsf{K})$ $=\sum_{(i,j)\in\mathsf{K}_{1}^{\star}}\mathsf{X}_{ij}-\sum_{(i,j)\in\mathsf{K}}\mathsf{X}_{ij}$ (105)
$=\sum_{(i,j)\in\mathsf{K}_{1}^{\star}}\mathbb{E}\mathsf{X}_{ij}-\sum_{(i,j)\in\mathsf{K}}\mathbb{E}\mathsf{X}_{ij}+\sum_{(i,j)\in\mathsf{K}_{1}^{\star}}[\mathsf{X}_{ij}-\mathbb{E}\mathsf{X}_{ij}]-\sum_{(i,j)\in\mathsf{K}}[\mathsf{X}_{ij}-\mathbb{E}\mathsf{X}_{ij}]$ (106)
@@
$=\lambda\cdot(k^{2}-\left|\mathsf{K}_{1}^{\star}\cap\mathsf{K}\right|)+\sum_{(i,j)\in\mathsf{K}_{1}^{\star}\setminus\mathsf{K}}[\mathsf{X}_{ij}-\lambda]-\sum_{(i,j)\in\mathsf{K}\setminus\mathsf{K}_{1}^{\star}}\mathsf{X}_{ij}$ (107)
$=\lambda\cdot(k^{2}-\left|\mathsf{K}_{1}^{\star}\cap\mathsf{K}\right|)+\mathsf{W}_{1}(\mathsf{K})+\mathsf{W}_{2}(\mathsf{K}),$ (108)
@@
where $\mathsf{W}_{1}(\mathsf{K})\triangleq\sum_{(i,j)\in\mathsf{K}_{1}^{\star}\setminus\mathsf{K}}[\mathsf{X}_{ij}-\lambda]$ and $\mathsf{W}_{2}(\mathsf{K})\triangleq-\sum_{(i,j)\in\mathsf{K}\setminus\mathsf{K}_{1}^{\star}}\mathsf{X}_{ij}$. Because $|\mathsf{K}|=|\mathsf{K}_{1}^{\star}|=k^{2}$, we have $|\mathsf{K}_{1}^{\star}\setminus\mathsf{K}|=|\mathsf{K}\setminus\mathsf{K}_{1}^{\star}|=k^{2}-\left|\mathsf{K}_{1}^{\star}\cap\mathsf{K}\right|$. Thus, both $\mathsf{W}_{1}(\mathsf{K})$ and $\mathsf{W}_{2}(\mathsf{K})$ are composed of the sum of $k^{2}-\left|\mathsf{K}_{1}^{\star}\cap\mathsf{K}\right|$ i.i.d. centered Gaussian random variables with unit variance. Accordingly, for $i=1,2$, and each fixed $\mathsf{K}$,
$\mathbb{P}\left(\mathsf{W}_{i}(\mathsf{K})\leq-\frac{\lambda}{2}(k^{2}-|\mathsf{K}_{1}^{\star}\cap\mathsf{K}|)\right)\leq\frac{1}{2}\exp\left[-\frac{1}{2}\lambda^{2}(k^{2}-|\mathsf{K}_{1}^{\star}\cap\mathsf{K}|)\right],$ (109)
and therefore, by the union bound and (108),
$\mathbb{P}\left(\mathsf{D}(\mathsf{K})\leq 0\right)\leq\exp\left[-\frac{1}{2}\lambda^{2}(k^{2}-|\mathsf{K}_{1}^{\star}\cap\mathsf{K}|)\right].$ (110)
Using (110) and the union bound once again, we get
@@
$\mathbb{P}\left(\hat{\mathsf{K}}_{1}(\mathsf{X})\neq\mathsf{K}_{1}^{\star}\right)$ $=\mathbb{P}\left[\bigcup_{\mathsf{K}\neq\mathsf{K}_{1}^{\star}}\mathsf{D}(\mathsf{K})\leq 0\right]$ (111)
$\leq\sum_{\mathsf{K}\neq\mathsf{K}_{1}^{\star}}\mathbb{P}\left(\mathsf{D}(\mathsf{K})\leq 0\right)$ (112)
$\leq\sum_{\mathsf{K}\neq\mathsf{K}_{1}^{\star}}\exp\left[-\frac{1}{2}\lambda^{2}(k^{2}-|\mathsf{K}_{1}^{\star}\cap\mathsf{K}|)\right]$ (113)
=$\stackrel{{\scriptstyle(a)}}{{=}}\sum_{\ell=0}^{k^{2}-k}\left|\mathsf{K}\in{\cal K}_{k,1,n}^{\mathsf{con}}:|\mathsf{K}_{1}^{\star}\cap\mathsf{K}|=\ell\right|e^{-\frac{1}{2}\lambda^{2}(k^{2}-\ell)}$ (114)
$\leq e^{-\frac{1}{2}\lambda^{2}k}\sum_{\ell=0}^{k^{2}-k}\left|\mathsf{K}\in{\cal K}_{k,1,n}^{\mathsf{con}}:|\mathsf{K}_{1}^{\star}\cap\mathsf{K}|=\ell\right|$ (115)
$\stackrel{{\scriptstyle(b)}}{{\leq}}n^{2}e^{-\frac{1}{2}\lambda^{2}k},$ (116)
@@
where in (a) we note that since $\mathsf{K}_{1}^{\star},\mathsf{K}\in{\cal K}_{k,1,n}^{\mathsf{con}}$ and $\mathsf{K}_{1}^{\star}\cap\mathsf{K}\neq\emptyset$, we must have that $|\mathsf{\mathsf{K}_{1}^{\star}}\cap\mathsf{K}|\leq k^{2}-k$, and $(b)$ is because $|{\cal K}_{k,1,n}^{\mathsf{con}}|\leq n^{2}$. Thus, we see that if $\lambda^{2}>\frac{(6+\epsilon)\log n}{k}$, then $\mathbb{P}(\hat{\mathsf{K}}_{1}(\mathsf{X})\neq\mathsf{K}_{1}^{\star})\leq n^{-(1+\epsilon/2)}$. Using the same steps above, it is clear that, $\mathbb{P}(\hat{\mathsf{K}}_{\ell}(\mathsf{X})\neq\mathsf{K}_{\ell}^{\star})\leq n^{-(1+\epsilon/2)}$, for any $2\leq\ell\leq m$, provided that $\lambda^{2}>\frac{(6+\epsilon)\log n}{k}$. Thus,
@@
$\mathbb{P}\left[\hat{\mathsf{K}}_{\mathsf{peel}}\neq\mathsf{K}^{\star}\right]=\mathbb{P}\left[\bigcup_{\ell=1}^{m}\hat{\mathsf{K}}_{\ell}\neq\mathsf{K}_{\ell}^{\star}\right]\leq\frac{m}{n^{(1+\epsilon/2)}}=n^{-\epsilon/2},$ (117)

### v2

@@ 1 Introduction
We use standard asymptotic notation. For two positive sequences $\{a_{n}\}$ and $\{b_{n}\}$, we write $a_{n}=O(b_{n})$ if $a_{n}\leq Cb_{n}$, for some absolute constant $C$ and for all $n$; $a_{n}=\Omega(b_{n})$, if $b_{n}=O(a_{n})$; $a_{n}=\Theta(b_{n})$, if $a_{n}=O(b_{n})$ and $a_{n}=\Omega(b_{n})$, $a_{n}=o(b_{n})$ or $b_{n}=\omega(a_{n})$, if $a_{n}/b_{n}\to 0$, as $n\to\infty$. Finally, for $a,b\in\mathbb{R}$, we let $a\vee b\triangleq\max\{a,b\}$ and $a\wedge b\triangleq\min\{a,b\}$. Throughout the paper, $\mathsf{C}$ refers to any constant independent of the parameters of the problem at hand and will be reused for different constants. The notation $\ll$ refers to polynomially less than in $n$, namely, $a_{n}\ll b_{n}$ if $\liminf_{n\to\infty}\log_{n}a_{n}<\liminf_{n\to\infty}\log_{n}b_{n}$, e.g., $n\ll n^{2}$, but $n\not\ll n\log_{2}n$. For $n\in\mathbb{N}$, we let $[n]=\{1,2,\dots,n\}$. For a subset $S\subseteq\mathbb{R}$, we let $\mathbbm{1}\{S\}$ denote the indicator function of the set $S$.
@@ 3.2 The recovery problem
Before we present this algorithm, we make a simplifying technical assumption on the possible set of planted submatrices, and then explain how this assumption can be removed. We assume that each pair of submatrices in the underlying planted submatrices $\mathsf{K}^{\star}$ are at least $k$ columns and rows far way. In other words, there are at least $k$ columns and $k$ rows separating any pair of submatrices in $\mathsf{K}^{\star}$. Similar assumptions are frequently taken when analyzing statistical models inspired by cryo-EM, see, for example [6]. We will refer to the above as the separation assumption.
@@
Consider the recovery problem in Definition 3, and let $\mathsf{C}$ be a universal constant. Then, we have the following set of bounds:
@@
    (ML Exact Recovery) Consider the ML estimator in (19). If
    $\liminf_{n\to\infty}\frac{\lambda}{\sqrt{\mathsf{C}k^{-1}\log n}}>1,$ (22)
@@
    (Peeling Exact Recovery) Consider the peeling estimator in (20), and assume that the separation assumption holds. Then, if
    $\liminf_{n\to\infty}\frac{\lambda}{\sqrt{\mathsf{C}k^{-1}\log n}}>1,$ (23)
    then exact recovery is possible.
  3. 3.
    (Peeling Correlated Recovery) Consider the peeling estimator in (20), and assume that the separation assumption holds. If
    $\liminf_{n\to\infty}\frac{\lambda}{\sqrt{\mathsf{C}k^{-2}\log n}}>1,$ (24)
@@ 4.5 Proof of Theorem 5
4.5.1 Exact recovery using the ML estimator
In this subsection, we analyze the ML estimator. Recall that,
$\hat{\mathsf{K}}_{\mathsf{ML}}(\mathsf{X})=\operatorname*{arg\,max}_{\mathsf{K}\in{\cal K}_{k,m,n}}{\cal S}(\mathsf{K}),$ (102)
where ${\cal S}(\mathsf{K})\triangleq\sum_{(i,j)\in\mathsf{K}}\mathsf{X}_{ij}$. We next prove the conditions for which $\hat{\mathsf{K}}_{\mathsf{ML}}=\mathsf{K}^{\star}$, with high probability, where $\mathsf{K}^{\star}$ are the $m$ planted submatrices. To prove the theorem, it suffices to show that ${\cal S}(\mathsf{K}^{\star})>{\cal S}(\mathsf{K})$, for all feasible $\mathsf{K}$ with $\mathsf{K}\neq\mathsf{K}^{\star}$. Let $\mathsf{D}(\mathsf{K})\triangleq{\cal S}(\mathsf{K}^{\star})-{\cal S}(\mathsf{K})$. Note that
$\mathsf{D}(\mathsf{K})$ $=\sum_{(i,j)\in\mathsf{K}^{\star}}\mathsf{X}_{ij}-\sum_{(i,j)\in\mathsf{K}}\mathsf{X}_{ij}$ (103)
$=\sum_{(i,j)\in\mathsf{K}^{\star}}\mathbb{E}\mathsf{X}_{ij}-\sum_{(i,j)\in\mathsf{K}}\mathbb{E}\mathsf{X}_{ij}+\sum_{(i,j)\in\mathsf{K}^{\star}}[\mathsf{X}_{ij}-\mathbb{E}\mathsf{X}_{ij}]-\sum_{(i,j)\in\mathsf{K}}[\mathsf{X}_{ij}-\mathbb{E}\mathsf{X}_{ij}]$ (104)
$=\lambda\cdot(mk^{2}-\left|\mathsf{K}^{\star}\cap\mathsf{K}\right|)+\sum_{(i,j)\in\mathsf{K}^{\star}\setminus\mathsf{K}}[\mathsf{X}_{ij}-\lambda]-\sum_{(i,j)\in\mathsf{K}\setminus\mathsf{K}^{\star}}\mathsf{X}_{ij}$ (105)
$=\lambda\cdot(mk^{2}-\left|\mathsf{K}^{\star}\cap\mathsf{K}\right|)+\mathsf{W}_{1}(\mathsf{K})+\mathsf{W}_{2}(\mathsf{K}),$ (106)
where $\mathsf{W}_{1}(\mathsf{K})\triangleq\sum_{(i,j)\in\mathsf{K}^{\star}\setminus\mathsf{K}}[\mathsf{X}_{ij}-\lambda]$ and $\mathsf{W}_{2}(\mathsf{K})\triangleq-\sum_{(i,j)\in\mathsf{K}\setminus\mathsf{K}^{\star}}\mathsf{X}_{ij}$. Because $|\mathsf{K}|=|\mathsf{K}^{\star}|=mk^{2}$, we have $|\mathsf{K}^{\star}\setminus\mathsf{K}|=|\mathsf{K}\setminus\mathsf{K}^{\star}|=mk^{2}-\left|\mathsf{K}^{\star}\cap\mathsf{K}\right|$. Thus, both $\mathsf{W}_{1}(\mathsf{K})$ and $\mathsf{W}_{2}(\mathsf{K})$ are composed of the sum of $mk^{2}-\left|\mathsf{K}^{\star}\cap\mathsf{K}\right|$ i.i.d. centered Gaussian random variables with unit variance. Accordingly, for $i=1,2$, and each fixed $\mathsf{K}$,
$\mathbb{P}\left(\mathsf{W}_{i}(\mathsf{K})\leq-\frac{\lambda}{2}(mk^{2}-|\mathsf{K}^{\star}\cap\mathsf{K}|)\right)\leq\frac{1}{2}\exp\left[-\frac{1}{2}\lambda^{2}(mk^{2}-|\mathsf{K}^{\star}\cap\mathsf{K}|)\right],$ (107)
and therefore, by the union bound and (106),
$\mathbb{P}\left(\mathsf{D}(\mathsf{K})\leq 0\right)\leq\exp\left[-\frac{1}{2}\lambda^{2}(mk^{2}-|\mathsf{K}^{\star}\cap\mathsf{K}|)\right].$ (108)
Using (108) and the union bound once again, we get
$\mathbb{P}\left(\hat{\mathsf{K}}_{\mathsf{ML}}(\mathsf{X})\neq\mathsf{K}^{\star}\right)$ $=\mathbb{P}\left[\bigcup_{\mathsf{K}\neq\mathsf{K}^{\star}}\mathsf{D}(\mathsf{K})\leq 0\right]$ (109)
$\leq\sum_{\mathsf{K}\neq\mathsf{K}^{\star}}\mathbb{P}\left(\mathsf{D}(\mathsf{K})\leq 0\right)$ (110)
$\leq\sum_{\mathsf{K}\neq\mathsf{K}^{\star}}\exp\left[-\frac{1}{2}\lambda^{2}(mk^{2}-|\mathsf{K}^{\star}\cap\mathsf{K}|)\right]$ (111)
$=\sum_{\ell=0}^{mk^{2}-k}\left|\mathsf{K}\in{\cal K}_{k,m,n}^{\mathsf{con}}:|\mathsf{K}^{\star}\cap\mathsf{K}|=\ell\right|e^{-\frac{1}{2}\lambda^{2}(mk^{2}-\ell)},$ (112)
where the last equality follows from the fact that since $\mathsf{K}^{\star},\mathsf{K}\in{\cal K}_{k,m,n}^{\mathsf{con}}$ and $\mathsf{K}^{\star}\cap\mathsf{K}\neq\emptyset$, we must have that $|\mathsf{\mathsf{K}^{\star}}\cap\mathsf{K}|\leq mk^{2}-k$. It can be shown that $\left|\mathsf{K}\in{\cal K}_{k,m,n}^{\mathsf{con}}:|\mathsf{K}^{\star}\cap\mathsf{K}|=\ell\right|\leq C\frac{(mk^{2}-\ell)^{2}}{k^{2}}n^{\frac{C^{\prime}(mk^{2}-\ell)}{k}}$, from some $C,C^{\prime}>0$, see, e.g., [23, Lemma 7]. Then,
$\mathbb{P}\left(\hat{\mathsf{K}}_{\mathsf{ML}}(\mathsf{X})\neq\mathsf{K}^{\star}\right)$ $\leq C\sum_{\ell=0}^{mk^{2}-k}\frac{(mk^{2}-\ell)^{2}}{k^{2}}n^{\frac{C^{\prime}(mk^{2}-\ell)}{k}}e^{-\frac{1}{2}\lambda^{2}(mk^{2}-\ell)}$ (113)
$=C\sum_{\ell=k}^{mk^{2}}\frac{\ell^{2}}{k^{2}}n^{\frac{C^{\prime}\ell}{k}}e^{-\frac{1}{2}\lambda^{2}\ell}$ (114)
$\leq C\sum_{\ell=k}^{mk^{2}}n^{4}n^{\frac{C^{\prime}\ell}{k}}e^{-\frac{1}{2}\lambda^{2}\ell}$ (115)
$=Cn^{4}\sum_{\ell=k}^{mk^{2}}e^{\frac{C^{\prime}\ell}{k}\log n-\frac{1}{2}\lambda^{2}\ell}$ (116)
$=Cn^{4}\sum_{\ell=k}^{mk^{2}}e^{-\ell\cdot\left(\frac{1}{2}\lambda^{2}-\frac{C^{\prime}}{k}\log n\right)}$ (117)
$\stackrel{{\scriptstyle(a)}}{{\leq}}Cn^{4}\sum_{\ell=k}^{mk^{2}}e^{-\frac{8\ell}{k}\log n}$ (118)
$\leq\frac{Cn^{4}mk^{2}}{n^{8}}=C\frac{mk^{2}}{n^{4}},$ (119)
where in (a) we have used the fact that $\lambda^{2}>\frac{(2C^{\prime}+16)\log n}{k}$. Thus, we get that $\mathbb{P}(\hat{\mathsf{K}}_{\mathsf{ML}}(\mathsf{X})\neq\mathsf{K}^{\star})$ converges to zero, as $n\to\infty$.
@@ 4.5.2 Exact recovery using the peeling estimator
where ${\cal S}(\mathsf{K})\triangleq\sum_{(i,j)\in\mathsf{K}}\mathsf{X}_{ij}$. We next prove the conditions for which $\hat{\mathsf{K}}_{1}(\mathsf{X})=\mathsf{K}^{\star}_{\ell}$, with high probability, for some $\ell\in[m]$, where $\mathsf{K}^{\star}=\cup_{\ell=1}^{m}\mathsf{K}_{\ell}^{\star}$ are the $m$ planted submatrices. To prove the theorem it suffices to show that ${\cal S}(\mathsf{K})>\max_{\ell\in[m]}{\cal S}(\mathsf{K}^{\star}_{\ell})$ is asymptotically small, for all feasible $\mathsf{K}$ with $\mathsf{K}\neq\mathsf{K}_{\ell}^{\star}$, for $\ell\in[m]$. Let $\mathsf{D}_{\ell}(\mathsf{K})\triangleq{\cal S}(\mathsf{K}^{\star}_{\ell})-{\cal S}(\mathsf{K})$. Note that
$\mathsf{D}_{\ell}(\mathsf{K})$ $=\sum_{(i,j)\in\mathsf{K}_{\ell}^{\star}}\mathsf{X}_{ij}-\sum_{(i,j)\in\mathsf{K}}\mathsf{X}_{ij}$ (121)
$=\sum_{(i,j)\in\mathsf{K}_{\ell}^{\star}}\mathbb{E}\mathsf{X}_{ij}-\sum_{(i,j)\in\mathsf{K}}\mathbb{E}\mathsf{X}_{ij}+\sum_{(i,j)\in\mathsf{K}_{\ell}^{\star}}[\mathsf{X}_{ij}-\mathbb{E}\mathsf{X}_{ij}]-\sum_{(i,j)\in\mathsf{K}}[\mathsf{X}_{ij}-\mathbb{E}\mathsf{X}_{ij}]$ (122)
@@
$=\lambda\cdot(k^{2}-\left|\mathsf{K}^{\star}\cap\mathsf{K}\right|)+\sum_{(i,j)\in\mathsf{K}_{\ell}^{\star}\setminus\mathsf{K}}[\mathsf{X}_{ij}-\lambda]-\sum_{(i,j)\in\mathsf{K}\setminus\mathsf{K}_{\ell}^{\star}}[\mathsf{X}_{ij}-\mathbb{E}\mathsf{X}_{ij}]$ (123)
$=\lambda\cdot(k^{2}-\left|\mathsf{K}^{\star}\cap\mathsf{K}\right|)+\mathsf{W}_{1}(\mathsf{K})+\mathsf{W}_{2}(\mathsf{K}),$ (124)
@@
where $\mathsf{W}_{1}(\mathsf{K})\triangleq\sum_{(i,j)\in\mathsf{K}_{\ell}^{\star}\setminus\mathsf{K}}[\mathsf{X}_{ij}-\lambda]$ and $\mathsf{W}_{2}(\mathsf{K})\triangleq-\sum_{(i,j)\in\mathsf{K}\setminus\mathsf{K}_{\ell}^{\star}}[\mathsf{X}_{ij}-\mathbb{E}\mathsf{X}_{ij}]$. Because $|\mathsf{K}|=|\mathsf{K}_{\ell}^{\star}|=k^{2}$, we have $|\mathsf{K}_{\ell}^{\star}\setminus\mathsf{K}|=|\mathsf{K}\setminus\mathsf{K}_{\ell}^{\star}|=k^{2}-\left|\mathsf{K}_{\ell}^{\star}\cap\mathsf{K}\right|$. Thus, both $\mathsf{W}_{1}(\mathsf{K})$ and $\mathsf{W}_{2}(\mathsf{K})$ are composed of sum of $k^{2}-\left|\mathsf{K}_{\ell}^{\star}\cap\mathsf{K}\right|$ i.i.d. centered Gaussian random variables with unit variance. Accordingly, for $i=1,2$, and each fixed $\mathsf{K}$,
$\mathbb{P}\left(\mathsf{W}_{i}(\mathsf{K})\leq-\frac{\lambda}{2}(k^{2}-|\mathsf{K}^{\star}\cap\mathsf{K}|)\right)$ $\leq\frac{1}{2}\exp\left[-\frac{\lambda^{2}}{8}\frac{(k^{2}-|\mathsf{K}^{\star}\cap\mathsf{K}|)^{2}}{k^{2}-\left|\mathsf{K}_{\ell}^{\star}\cap\mathsf{K}\right|}\right].$ (125)
Therefore, by the union bound and (124),
$\mathbb{P}\left(\mathsf{D}_{\ell}(\mathsf{K})\leq 0\right)\leq\exp\left[-\frac{\lambda^{2}}{8}\frac{(k^{2}-|\mathsf{K}^{\star}\cap\mathsf{K}|)^{2}}{k^{2}-\left|\mathsf{K}_{\ell}^{\star}\cap\mathsf{K}\right|}\right].$ (126)
Note that due to the separation assumption, it must be the case that either $|\mathsf{K}^{\star}\cap\mathsf{K}|=|\mathsf{K}_{j}^{\star}\cap\mathsf{K}|\neq 0$, for some $j\in[m]$, or $|\mathsf{K}^{\star}\cap\mathsf{K}|=0$. In the later case, we have
$\mathbb{P}\left(\mathsf{D}_{\ell}(\mathsf{K})\leq 0\right)\leq\exp\left[-\frac{\lambda^{2}k^{2}}{8}\right],$ (127)
while in the former the exists a unique $j\in[m]$, such that,
$\min_{\ell\in[m]}\mathbb{P}\left(\mathsf{D}_{\ell}(\mathsf{K})\leq 0\right)$ $\leq\min_{\ell\in[m]}\exp\left[-\frac{\lambda^{2}}{8}\frac{(k^{2}-|\mathsf{K}_{j}^{\star}\cap\mathsf{K}|)^{2}}{k^{2}-\left|\mathsf{K}_{\ell}^{\star}\cap\mathsf{K}\right|}\right]$ (128)
$\leq\exp\left[-\frac{\lambda^{2}}{8}(k^{2}-|\mathsf{K}_{j}^{\star}\cap\mathsf{K}|)\right]$ (129)
$\leq\exp\left[-\frac{\lambda^{2}k}{8}\right],$ (130)
where the third inequity is since $\mathsf{K}_{j}^{\star},\mathsf{K}\in{\cal K}_{k,1,n}^{\mathsf{con}}$ and $\mathsf{K}_{j}^{\star}\cap\mathsf{K}\neq\emptyset$, we must have that $|\mathsf{\mathsf{K}_{j}^{\star}}\cap\mathsf{K}|\leq k^{2}-k$. Accordingly, using (126) and the union bound once again, we get
@@
$\mathbb{P}\left(\hat{\mathsf{K}}_{1}(\mathsf{X})\neq\mathsf{K}_{\ell}^{\star}\;\text{for some }\ell\in[m]\right)$ $=\mathbb{P}\left[\bigcup_{\mathsf{K}\neq(\mathsf{K}_{1}^{\star},\ldots,\mathsf{K}_{m}^{\star})}\left\{{\cal S}(\mathsf{K})>\max_{\ell\in[m]}{\cal S}(\mathsf{K}^{\star}_{\ell})\right\}\right]$ (131)
$=\mathbb{P}\left[\bigcup_{\mathsf{K}\neq(\mathsf{K}_{1}^{\star},\ldots,\mathsf{K}_{m}^{\star})}\left\{\mathsf{D}_{1}(\mathsf{K})\leq 0,\ldots,\mathsf{D}_{m}(\mathsf{K})\leq 0\right\}\right]$ (132)
$\leq\sum_{\mathsf{K}\neq(\mathsf{K}_{1}^{\star},\ldots,\mathsf{K}_{m}^{\star})}\min_{\ell\in[m]}\mathbb{P}\left(\mathsf{D}_{\ell}(\mathsf{K})\leq 0\right)$ (133)
$\leq\sum_{\mathsf{K}\neq(\mathsf{K}_{1}^{\star},\ldots,\mathsf{K}_{m}^{\star})}\exp\left[-\frac{\lambda^{2}k}{8}\right]$ (134)
$\leq n^{2}e^{-\frac{1}{8}\lambda^{2}k},$ (135)
@@
where the last inequality is because $|{\cal K}_{k,1,n}^{\mathsf{con}}|\leq n^{2}$. Thus, we see that if $\lambda^{2}>\frac{(24+\epsilon)\log n}{k}$, then $\mathbb{P}\left(\hat{\mathsf{K}}_{1}(\mathsf{X})\neq\mathsf{K}_{\ell}^{\star}\;\text{for some }\ell\in[m]\right)\leq n^{-(1+\epsilon/8)}$. Using the same steps above, it is clear that, $\mathbb{P}(\hat{\mathsf{K}}_{\ell}(\mathsf{X})\neq\mathsf{K}_{\ell}^{\star})\leq n^{-(1+\epsilon/8)}$, for any $2\leq\ell\leq m$, provided that $\lambda^{2}>\frac{(24+\epsilon)\log n}{k}$. Thus,
@@
$\mathbb{P}\left[\hat{\mathsf{K}}_{\mathsf{peel}}\neq\mathsf{K}^{\star}\right]=\mathbb{P}\left[\bigcup_{\ell=1}^{m}\hat{\mathsf{K}}_{\ell}\neq\mathsf{K}_{\ell}^{\star}\right]\leq\frac{m}{n^{(1+\epsilon/8)}}=n^{-\epsilon/8},$ (136)